Find the equation of a quadratic polynomial passing through the following points: (-1,2) (-5,2) and (6,8)?
Answer by venugopalramana(3286)

Find the equation of a quadratic polynomial passing through the following points: A=(-1,2).....B= (-5,2) and C=(6,8)?
EQN. OF AB IS
Y-2 = {(2-2)/(-5+1)}(X+1)=0
Y-2=0...............................1
EQN.OF BC IS
Y-2={(8-2)/(6+5)}(X+5)=6(X+5)/11
11Y-22=6X+30
11Y-6X-30-22=0
11Y-6X-52=0...................2
HENCE THE COMMON EQN.REPRESENTING BOTH THE EQNS. IS
(Y-2)(11Y-6X-52)=0
11Y^2-6XY-52Y-22Y-12X+104=0
11Y^2-6XY-74Y-12X+104=0